# PointSeg: A Training-Free Paradigm for 3D Scene Segmentation via Foundation Models

### v4

@@ 4.1 Setup
We compare our approach with both training-based and training-free baselines. For training-based comparison, we select the state-of-the-art transformer-based method PointSAM (Zhou et al., 2024). For training-free methods, we compare against the 2D-to-3D lifting methods (Yang et al., 2023b; Xu et al., 2023) and the 3D-to-2D projection methods (Xu et al., 2023) respectively. For Implementation, we apply the V-DETR (Shen et al., 2024) and VirConv (Wu et al., 2023b) as the indoor and outdoor 3D detector.
@@ 4.2 Main Results
Comparisons with the state-of-the-art Methods.
We compare the segmentation results on ScanNet, ScanNet++, and KITTI-360 datasets, covering both indoor and outdoor scenes. As shown in Table 1, comparing to previous training-free methods, our PointSeg obtains 16.3 mAP, 18 AP50, and 16.4 AP25 performance gains on ScanNet. On the more challenging indoor dataset ScanNet++, our method still obtains 14.9 mAP, 15.4 AP50, and 15.6 AP25 improvements. Furthermore, when evaluating the performance of our PointSeg on outdoor KITTI-360, our method still surpasses corresponding zero-shot method by 15 mAP, 16.1 AP50, and 16.9 AP25, respectively. In this regard, our PointSeg demonstrates superior generalization ability to complex 3D scenarios.
Notably, when compared to previous training-based methods, PointSeg outperforms PointSAM (Zhou et al., 2024) by 5.6-8, 7.2-11.1, 7.9-10.3 in terms of mAP, AP50, and AP25 across various datasets. This further demonstrates the robustness and effectiveness of our approach in 3D segmentation task.
Qualitative Results.
The representative quantitative segmentation results of our proposed PointSeg on the three datasets are shown in Figure 1 and Figure A1. We also present the quantitative results of the state-of-the-art training-based methods PointSAM (Zhou et al., 2024) and training-free method SAMPro3D (Xu et al., 2023). We can observe that PointSAM and SAMPro3D often mistakenly segment an object into two different objects and exhibit poor performance in segmenting relative objects lacked spatial structure. Our PointSeg can handle complex scenes and is capable of generating clean segments on objects of small size, further underscoring the effectiveness of our approach.
@@ 4.3 Ablation Study
In this section, we conduct extensive ablation studies on ScanNet to show the effectiveness of each component and design. Unless otherwise specified, SAM (Kirillov et al., 2023) is used as 2D segmentation foundation model for ablation studies by default.
@@
Different Foundation Models. i) Apart the basic segmentation foundation model SAM 2 (Ravi et al., 2024), we also integrate different segmentation foundation model, i.e., SAM (Kirillov et al., 2023), MobileSAM (Zhang et al., 2023a), FastSAM (Zhao et al., 2023), and EfficientSAM (Xiong et al., 2023), into our framework. As shown in Table 2, PointSeg demonstrates consistent performance improvements among different datasets, where the original SAM 2-based result performs best. This is consistent with the relative results of these methods in 2D segmentation. ii) In Table 4.1, we change the model for localization of point prompts. The results demonstrate that more accurate points for 3D prompts can indeed contribute to the performance gain. iii) In Table 4.1, the improvement of different 3D detectors will also bring about the improvement of the performance of our method. These results suggest that our framework can serve as a foundational structure, capable of integrating a variety of fundamental models. And the enhancements observed in these models can be smoothly translated to 3D space, thereby augmenting the overall performance.
@@
Main Components. Further, we explore the effects of bidirectional matching based prompts generation (BMP), iterative post-refinement (IPR) and affinity-aware merging (AM). We illustrate the importance of different components by removing some parts and keeping all the others unchanged. The baseline setting is to use points and boxes independently as 3D prompts. And the masks from the 2D segmentation model decoder are used to merge the final 3D masks, without any post-refinement. The merging strategy is same as the adjacent frame merging in (Yang et al., 2023b). The results of components ablations are shown in Table 4.3. We observe that using the iterative post-refinement strategy leads to a noticeable increase in performance, which demonstrates the necessity of refining the initial 2D masks. The performance degradation caused by the absence of bidirectional matching proves that the constraints between the point and box prompts can indeed help to generate the accurate point-box pairs. And the performance drop without the affinity-aware merging shows that the affinity score is indeed useful to link the point and the masks.
BMP IPR AM mAP AP50 AP25 - - - 16.2 40.6 60.3 - - ✓ 28.6 50.9 69.7 - ✓ - 32.5 55.2 74.3 ✓ - - 29.5 51.9 71.2 ✓ ✓ - 33.9 57.8 76.5 ✓ - ✓ 31.7 53.3 75.1 - ✓ ✓ 34.6 58.2 77.5 ✓ ✓ ✓ 36.3 60.2 79.3
Strategy mAP AP50 AP25 Iter 0 30.3 54.8 73.8 1 31.9 55.6 74.1 2 32.6 56.8 75.7 3 34.7 58.5 77.8 4 32.3 56.9 75.5 5 30.4 54.1 73.2 Adaptive 36.3 60.2 79.3
Strategy mAP AP50 AP25 no 22.2 49.6 67.3 forward 30.9 55.5 73.5 reverse 32.6 56.7 75.5 bidirectional 36.3 60.2 79.3
Merging mAP AP50 AP25 BM 27.6 53.5 73.7 PM 31.5 55.7 76.2 IDM 32.1 56.2 76.8 AM 36.3 60.2 79.3
Strategy mAP AP50 AP25 PointClip only 30.2 51.2 72.7 3D detector only 30.1 51.3 72.6 combine(w/o matching) 34.6 58.2 77.5 combine(w/ matchfing) 36.3 60.2 79.3
@@
Iterative Post-refinement. As mentioned before, when performing the post-refinement during the initial 2D mask generation, we have tried different strategies of fixed numbers method and adaptive iteration method. As shown in Table 4.3, the six rows in the middle represent the method to use a fixed number of iterations and the last row is the adaptive iteration. We can observe that as the number of iterations increases, the corresponding AP value also becomes higher compared to no iteration, which shows that the obtained mask is also more accurate. In the method with a fixed number of iterations, the results reach the highest in the third iteration, but are still lower than those in the adaptive iteration method. This largely illustrates the effectiveness of the iterative post-processing method in generating more accurate masks, and also reveals that the adaptive iteration is more beneficial.
Ratio mAP AP50 AP25 1 33.6 57.5 76.6 3 35.2 59.3 78.7 5 36.3 60.2 79.3 8 35.8 59.7 78.1 10 34.6 58.3 77.9 15 32.2 56.6 75.5
@@
Different Matching Strategy. To validate the effect of different matching method, we explore the effects of the forward matching and the reverse matching of the proposed bidirectional matching, as shown in Table 4.3. Without the guidance from the respective point and box, the naive prompts contain many invalid points and regions, which provide negative prompts for the following segmentation models. Our bidirectional matching improves the performance of forward and reverse matching by 5.4 and 3.7 , which show the effectiveness of the proposed bidirectional matching strategy.
@@
module bidirectional matching iterative refinement affinity-aware merging FPS 1.53 1.05 1.96
Affinity-aware Merging. In the final mask merging stage, we have also tried other merging algorithm. As shown in Table 4.3, we compare our affinity-aware merging (AM) with (a) bidirectional merging (BM) from (Yang et al., 2023b), (b) pure merging (PM) without affinity scores, which is simplified from our approach and (c) prompt ID based merging (IDM) from (Xu et al., 2023). With other inferior merging method, the performance drops dramatically which shows the superiority of our proposed affinity-aware merging algorithm in solving the mask merging problems in 3D scene.
@@
3D Point Models. As shown in Table 4.3, the missing of PointLLM/3D detector causes the performance drop and only the matching of combination performs best.
Mask Change Ratio. In the iterative post-refinement module, we set a mask change ratio threshold as a condition for stopping the iteration. Here, We show the effect of different ratios on the results. As shown in Table 10, the overall results perform best when mask change ratio is set as 5.
@@
Inference Speed. As shown in Table 11, we test the running efficiency on NVIDIA V100 GPU and the detailed FPS of each module. The FLOPs of our method is 4.2G.
@@ 5 Conclusion
Appendix
Appendix A Datasets and Metrics.
To validate the effectiveness of our proposed PointSeg, we conduct extensive experiments on three popular public benchmarks: ScanNet (Dai et al., 2017), ScanNet++ (Yeshwanth et al., 2023), and KITTI-360 (Liao et al., 2022). ScanNet provides RGBD images and 3D meshes of 1613 indoor scenes. ScanNet++ is a recently released indoor dataset with more detailed segmentation masks, serving as a more challenging benchmark for 3D scenarios. It contains 280 indoor scenes with high-fidelity geometry and high-resolution RGB images. KITTI-360 is a substantial outdoor dataset that includes 300 suburban scenes, which comprises 320k images and 100k laser scans. We evaluate ours segmentation performance with the widely-used Average Precision (AP) score. Following (Schult et al., 2023; Dai et al., 2017), we report AP with thresholds of 50 and 25 (denoted as AP50 and AP25) as well as AP averaged with IoU thresholds from 50 to 95 with a step size of 5 (mAP).

### v5

@@ 4.1 Setup
Method Type ScanNet ScanNet++ KITTI-360 mAP AP50 AP25 mAP AP50 AP25 mAP AP50 AP25 With Training SAM-graph [11] specialist model 15.1 33.3 59.1 12.9 25.3 43.6 14.7 28.0 43.2 Mask3D [42] specialist model 26.9 44.4 57.5 8.8 15.0 22.3 0.1 0.4 4.2 OpenDAS [62] specialist model 28.3 49.6 66.2 19.2 35.5 52.6 20.1 32.4 52.2 PointSAM [70] specialist model 32.9 56.4 72.5 25.8 38.0 59.3 25.1 38.4 56.2 Training-free SAM3D [59] specialist model 13.7 29.7 54.5 8.3 17.5 33.7 6.3 16.0 35.6 SAI3D [64] specialist model 18.8 42.5 62.3 17.1 31.1 49.5 16.5 30.2 48.6 SAMPro3D [53] specialist model 22.2 45.6 65.7 18.9 33.7 51.6 17.3 31.1 49.6 PointSeg (Ours) generalist model 38.5 63.6 82.1 33.8 49.1 67.2 32.3 47.2 66.5
Method ScanNet ScanNet++ KITTI-360 mAP AP50 AP25 mAP AP50 AP25 mAP AP50 AP25 + SAM 2 [40] 38.5 63.6 82.1 33.8 49.1 67.2 32.3 47.2 66.5 + SAM [24] 36.3 60.2 79.3 31.2 46.5 64.8 29.9 44.5 63.3 + MobileSAM [65] 26.2 49.8 68.3 19.6 36.4 55.2 20.6 34.6 53.3 + FastSAM [69] 26.9 50.8 69.1 20.5 37.7 56.5 21.2 35.8 54.5 + EfficientSAM [52] 33.5 57.2 75.8 27.8 43.6 62.4 27.5 41.7 60.6
@@
We compare our approach with both training-based and training-free baselines. For training-based comparison, we select the state-of-the-art transformer-based method PointSAM [70]. For training-free methods, we compare against the 2D-to-3D lifting methods [59, 53] and the 3D-to-2D projection methods [53] respectively. For Implementation, we apply the V-DETR [43] and VirConv [51] as the indoor and outdoor 3D detector.
@@ 4.2 Main Results
Comparisons with the state-of-the-art Methods. We compare the segmentation results on ScanNet, ScanNet++, and KITTI-360 datasets, covering both indoor and outdoor scenes. As shown in Tab. 1, comparing to previous training-free methods, our PointSeg obtains 16.3 mAP, 18 AP50, and 16.4 AP25 performance gains on ScanNet. On the more challenging indoor dataset ScanNet++, our method still obtains 14.9 mAP, 15.4 AP50, and 15.6 AP25 improvements. Furthermore, when evaluating the performance of our PointSeg on outdoor KITTI-360, our method still surpasses corresponding zero-shot method by 15 mAP, 16.1 AP50, and 16.9 AP25, respectively. In this regard, our PointSeg demonstrates superior generalization ability to complex 3D scenarios.
Notably, when compared to previous training-based methods, PointSeg outperforms PointSAM [70] by 5.6-8, 7.2-11.1, 7.9-10.3 in terms of mAP, AP50, and AP25 across various datasets. This further demonstrates the robustness and effectiveness of our approach in 3D segmentation task.
Qualitative Results. The representative quantitative segmentation results of our proposed PointSeg on the three datasets are shown in Fig. 1 . We also present the quantitative results of the state-of-the-art training-based methods PointSAM [70] and training-free method SAMPro3D [53]. We can observe that PointSAM and SAMPro3D often mistakenly segment an object into two different objects and exhibit poor performance in segmenting relative objects lacked spatial structure. Our PointSeg can handle complex scenes and is capable of generating clean segments on objects of small size, further underscoring the effectiveness of our approach.
@@ 4.3 Ablation Study
In this section, we conduct extensive ablation studies on ScanNet to show the effectiveness of each component and design. Unless otherwise specified, SAM [24] is used as 2D segmentation foundation model for ablation studies by default.
@@
Different Foundation Models. i) Apart the basic segmentation foundation model SAM 2 [40], we also integrate different segmentation foundation model, i.e., SAM [24], MobileSAM [65], FastSAM [69], and EfficientSAM [52], into our framework. As shown in Tab. 2, PointSeg demonstrates consistent performance improvements among different datasets, where the original SAM 2-based result performs best. This is consistent with the relative results of these methods in 2D segmentation. ii) In Tab. 3, we change the model for localization of point prompts. The results demonstrate that more accurate points for 3D prompts can indeed contribute to the performance gain. iii) In Tab. 4, the improvement of different 3D detectors will also bring about the improvement of the performance of our method. These results suggest that our framework can serve as a foundational structure, capable of integrating a variety of fundamental models. And the enhancements observed in these models can be smoothly translated to 3D space, thereby augmenting the overall performance.
BMP | IPR | AM | mAP | AP50 | AP25 |
- | - | - | 16.2 | 40.6 | 60.3 |
- | - | ✓ | 28.6 | 50.9 | 69.7 |
- | ✓ | - | 32.5 | 55.2 | 74.3 |
✓ | - | - | 29.5 | 51.9 | 71.2 |
✓ | ✓ | - | 33.9 | 57.8 | 76.5 |
✓ | - | ✓ | 31.7 | 53.3 | 75.1 |
- | ✓ | ✓ | 34.6 | 58.2 | 77.5 |
✓ | ✓ | ✓ | 36.3 | 60.2 | 79.3 |
@@
Main Components. Further, we explore the effects of bidirectional matching based prompts generation (BMP), iterative post-refinement (IPR) and affinity-aware merging (AM). We illustrate the importance of different components by removing some parts and keeping all the others unchanged. The baseline setting is to use points and boxes independently as 3D prompts. And the masks from the 2D segmentation model decoder are used to merge the final 3D masks, without any post-refinement. The merging strategy is same as the adjacent frame merging in [59]. The results of components ablations are shown in Tab. 5. We observe that using the iterative post-refinement strategy leads to a noticeable increase in performance, which demonstrates the necessity of refining the initial 2D masks. The performance degradation caused by the absence of bidirectional matching proves that the constraints between the point and box prompts can indeed help to generate the accurate point-box pairs. And the performance drop without the affinity-aware merging shows that the affinity score is indeed useful to link the point and the masks.
Strategy | mAP | AP50 | AP25 | |
Iter | 0 | 30.3 | 54.8 | 73.8 |
1 | 31.9 | 55.6 | 74.1 | |
2 | 32.6 | 56.8 | 75.7 | |
3 | 34.7 | 58.5 | 77.8 | |
4 | 32.3 | 56.9 | 75.5 | |
5 | 30.4 | 54.1 | 73.2 | |
Adaptive | 36.3 | 60.2 | 79.3 |
@@
Iterative Post-refinement. As mentioned before, when performing the post-refinement during the initial 2D mask generation, we have tried different strategies of fixed numbers method and adaptive iteration method. As shown in Tab. 6, the six rows in the middle represent the method to use a fixed number of iterations and the last row is the adaptive iteration. We can observe that as the number of iterations increases, the corresponding AP value also becomes higher compared to no iteration, which shows that the obtained mask is also more accurate. In the method with a fixed number of iterations, the results reach the highest in the third iteration, but are still lower than those in the adaptive iteration method. This largely illustrates the effectiveness of the iterative post-processing method in generating more accurate masks, and also reveals that the adaptive iteration is more beneficial.
Strategy | mAP | AP50 | AP25 |
no | 22.2 | 49.6 | 67.3 |
forward | 30.9 | 55.5 | 73.5 |
reverse | 32.6 | 56.7 | 75.5 |
bidirectional | 36.3 | 60.2 | 79.3 |
@@
Different Matching Strategy. To validate the effect of different matching method, we explore the effects of the forward matching and the reverse matching of the proposed bidirectional matching, as shown in Tab. 7. Without the guidance from the respective point and box, the naive prompts contain many invalid points and regions, which provide negative prompts for the following segmentation models. Our bidirectional matching improves the performance of forward and reverse matching by 5.4 and 3.7 , which show the effectiveness of the proposed bidirectional matching strategy.
Strategy | mAP | AP50 | AP25 |
PointClip only | 30.2 | 51.2 | 72.7 |
3D detector only | 30.1 | 51.3 | 72.6 |
combine(w/o matching) | 34.6 | 58.2 | 77.5 |
combine(w/ matchfing) | 36.3 | 60.2 | 79.3 |
@@
module | bidirectional matching | iterative refinement | affinity-aware merging |
FPS | 1.53 | 1.05 | 1.96 |
@@
3D Point Models. As shown in Tab. 8, the missing of PointLLM/3D detector causes the performance drop and only the matching of combination performs best.
@@
Inference Speed. As shown in Tab. 9, we test the running efficiency on NVIDIA V100 GPU and the detailed FPS of each module. The FLOPs of our method is 4.2G.
@@ 5 Conclusion
Supplementary Material
The supplementary material presents the following sections to strengthen the main manuscript:
Datasets and Metrics.
More Ablation Studies.
A1 Datasets and Metrics
To validate the effectiveness of our proposed PointSeg, we conduct extensive experiments on three popular public benchmarks: ScanNet [4], ScanNet++ [61], and KITTI-360 [33]. ScanNet provides RGBD images and 3D meshes of 1613 indoor scenes. ScanNet++ is a recently released indoor dataset with more detailed segmentation masks, serving as a more challenging benchmark for 3D scenarios. It contains 280 indoor scenes with high-fidelity geometry and high-resolution RGB images. KITTI-360 is a substantial outdoor dataset that includes 300 suburban scenes, which comprises 320k images and 100k laser scans. We evaluate ours segmentation performance with the widely-used Average Precision (AP) score. Following [42, 4], we report AP with thresholds of 50 and 25 (denoted as AP50 and AP25) as well as AP averaged with IoU thresholds from 50 to 95 with a step size of 5 (mAP).
A2 More Ablation Studies
Merging | mAP | AP50 | AP25 |
BM | 27.6 | 53.5 | 73.7 |
PM | 31.5 | 55.7 | 76.2 |
IDM | 32.1 | 56.2 | 76.8 |
AM | 36.3 | 60.2 | 79.3 |
Affinity-aware Merging. In the final mask merging stage, we have also tried other merging algorithm. As shown in Tab. A1, we compare our affinity-aware merging (AM) with (a) bidirectional merging (BM) from [59], (b) pure merging (PM) without affinity scores, which is simplified from our approach and (c) prompt ID based merging (IDM) from [53]. With other inferior merging method, the performance drops dramatically which shows the superiority of our proposed affinity-aware merging algorithm in solving the mask merging problems in 3D scene.
Ratio | mAP | AP50 | AP25 |
1 | 33.6 | 57.5 | 76.6 |
3 | 35.2 | 59.3 | 78.7 |
5 | 36.3 | 60.2 | 79.3 |
8 | 35.8 | 59.7 | 78.1 |
10 | 34.6 | 58.3 | 77.9 |
15 | 32.2 | 56.6 | 75.5 |
Mask Change Ratio. In the iterative post-refinement module, we set a mask change ratio threshold as a condition for stopping the iteration. Here, We show the effect of different ratios on the results. As shown in Tab. A2, the overall results perform best when mask change ratio is set as 5.